# TorchTitan: One-stop PyTorch native solution for production ready LLM pre-training

### v2

@@ 2.2.1 Navigating compute-memory trade-offs using activation checkpointing
Activation checkpointing (AC) (Chen et al., 2016) and selective activation checkpointing (SAC) (Korthikanti et al., 2023) are standard training techniques to reduce peak GPU memory usage, by trading activation recomputation during the backward pass for memory savings. It is often needed even after applying multi-dimensional parallelisms.
@@
Within a TransformerBlock, full AC works by recomputing all activation tensors needed during the backward pass, whereas op-level SAC saves the results from computation-intensive PyTorch operations and only recomputes others. Layer-level SAC works in similar fashion as full AC, but the wrapping is applied to every $x$ TransformerBlock (where $x$ is specified by the user) to implement configurable trade-offs between memory and recompute. (Details are in Appendix  9.5.)
@@
In TorchTitan, we use regional compilation, which applies torch.compile to each individual TransformerBlock in the Transformer model. This has two main benefits: (1) we get a full graph (without graph breaks) for each region, compatible with FSDP2 and TP (and more generally torch.Tensor subclasses such as DTensor) and other PyTorch distributed training techniques; (2) since the Llama model stacks identical TransformerBlock layers one after another, torch.compile can identify the same structure being repeatedly compiled and only compile once, thus greatly reducing compilation time.
@@ 2.2.3 Asynchronous Tensor Parallel to maximally overlap communication
By default TP incurs blocking communications before/after the sharded computations, causing computation resources to not be effectively utilized. Asynchronous TP (AsyncTP) (Wang et al., 2022) achieves computation-communication overlap by fractionalizing the TP matrix multiplications within the attention and feed-forward modules into smaller chunks, and overlapping communication collectives in between each section. The overlap is achieved by a micro-pipelining optimization, where results are being communicated at the same time that the other chunks of the matmul are being computed.
PyTorch AsyncTP is based on a SymmetricMemory abstraction, which creates intra-node buffers to write faster communication collectives (Wang et al., 2024). This is done by allocating an shared memory buffer on each GPU in order to provide direct P2P access.
@@ 2.2.4 Boosting throughput with mixed precision training and Float8 support
Mixed precision training (Micikevicius et al., 2018) provides both memory and computational savings while ensuring training stability. FSDP2 has built-in support for mixed precision training with basic torch.dtype. This covers the popular usage of performing FSDP all-gather and computation in a low precision (e.g. torch.bfloat16), and perform lossless FSDP reduce-scatter (gradient) in high precision (e.g. torch.float32) for better numerical results. See Appendix  9.7 for usage details.
TorchTitan also supports more advanced mixed precision training with Float8 (a derived data type) on newer hardware like H100, with substantial performance gains (reported in Section 3.2). The Float8 feature from torchao.float8 supports multiple per-tensor scaling strategies, including dynamic, delayed, and static (see Micikevicius et al. (2022); Vasiliy Kuznetsov (2024), Section 4.3 for details), while being composable with other key PyTorch-native systems such as autograd, torch.compile, FSDP2 and TP (with Float8 all-gather capability (Feng et al., 2024)).
@@ 2.3.1 Scalable and efficient Distributed Checkpointing
Checkpoints are crucial in training large language models for two reasons: they facilitate model reuse in applications like inference and evaluation, and they provide a recovery mechanism in case of failures. An optimal checkpointing workflow should ensure ease of reuse across different parallelisms and maintain high performance without slowing down training. There are two typical checkpointing methods. The first aggregates the state (model parameters and optimizer states) into an unsharded version that is parallelism-agnostic, facilitating easy reuse but requiring expensive communication. The second method has each trainer save its local sharded state, which speeds up the process but complicates reuse due to embedded parallelism information.
DCP addresses these challenges using DTensor, which encapsulates both global and local tensor information independently of parallelism. DCP converts this information into an internal format for storage. During loading, DCP matches the stored shards with the current DTensor-based model parameters and optimizer states, fetching the necessary shard from storage. TorchTitan, which utilizes all native PyTorch parallelisms, effectively uses DCP to balance efficiency and usability. Furthermore, DCP enhances efficiency through asynchronous checkpointing by processing storage persistence in a separate thread, allowing this operation to overlap with subsequent training iterations. TorchTitan utilizes DCP’s asynchronous checkpointing to reduce the checkpointing overhead by 5-15x compared to synchronous distributed checkpointing for the Llama 3.1 8B model (Zhang et al., 2024; Huang et al., 2024).
@@ 2.3.2 Flight Recorder to debug job crashes
A common failure mode when developing parallelism code, or when running at large scale, is to observe a NCCL collective timeout and then the need to figure out the root cause. Since communication kernels are typically asynchronous from the perspective of the CPU, by the time something times out, it can be very hard to pinpoint which operation failed and why. PyTorch provides a Flight Recorder for NCCL collectives to help resolve this dilemma. It records the start and end time (on GPU) as well as the enqueue time (on CPU) for every collective or p2p operation. Additionally, it logs metadata such as which process group was used, who the source rank is (and destination, for p2p), tensor sizes, and stack traces.
We find the data contained in the Flight Recorder helps debug collective hangs and p2p hangs caused by bugs in parallelism code. For PP, there may be schedule bugs that lead to hangs, due to a missing or improperly ordered send or recv operation. Analysis based on the Flight Recorder data can pinpoint the latest send or recv that has completed on the GPU. For FSDP or TP, it is possible to determine whether one or more ranks has not called into a collective, perhaps due to a bug in PP scheduling or faulty logic in TP.
@@ 3 Experimentation
In this section, we demonstrate the effectiveness of elastic distributed training using TorchTitan, via experiments on Llama 3.1 8B, 70B, and 405B, from 1D parallelism to 3D parallelism (respectively), at the scale from 8 GPUs to 512 GPUs. We also share the knowledge and experience gained through TorchTitan experimentation. A walkthrough of the codebase on how we apply (up to) 3D parallelism can be found in Appendix 8.
@@ 3.2 Performance
To showcase the elasticity and scalability of TorchTitan, we experiment on a wide range of GPU scales (from 8 to 512), as the underlying model size increases (8B, 70B, and 405B) with a varying number of parallelism dimensions (1D, 2D, and 3D, respectively). To demonstrate the effectiveness of the optimization techniques introduced in Section 2.2, we show how training throughput improves when adding each individual technique on appropriate baselines. In particular, when training on a higher dimensional parallelism with new features, the baseline is always updated to include all previous techniques.
@@
Table 1: 1D parallelism (FSDP) on Llama 3.1 8B model, 8 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 16.
@@
Table 2: 1D parallelism (FSDP) on Llama 3.1 8B model, 128 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 256.
@@
Table 3: 2D parallelism (FSDP + TP) + torch.compile + Float8 on Llama 3.1 70B model, 256 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 32, TP degree 8. Local batch size 16, global batch size 512.
@@
Table 4: 3D parallelism (FSDP + TP + PP) + torch.compile + Float8 + AsyncTP on Llama 3.1 405B model, 512 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 4, TP degree 8, PP degree 16. Local batch size 32, global batch size 128.
@@
3.3 Scaling with TorchTitan 3D parallelism
The LLM scaling law imposes challenges due to increasingly larger model size and huge amount of data, which requires applying parallelism strategies on massive number of GPUs. TorchTitan provides the ability to compose different parallelisms to efficiently scale model training to thousands of GPUs. This section discusses the observations and motivations to apply TorchTitan 3D parallelism when training LLMs at large scale. Please note that there could be many 3D parallelism combinations, but we choose to only discuss one combination in this paper, which could be summarized as the following diagram:
Figure 2: Scaling with 3D Parallelism
@@ 3.3.1 Scaling with FSDP
FSDP (ZeRO) is a generalized technique which can be applied to any model architecture, making it a good choice for the first or the only degree of parallelism. As long as the FSDP communication is faster than the corresponding computation (which is the case for LLMs trained on up to hundreds of, say 512, GPUs), and there is no need to reduce (effective) per-GPU batch size to be below 1 (for reasons mentioned below in the TP section), 1D FSDP should be sufficient.
Existing ring-based implementations of NCCL collectives (all-gather, reduce-scatter) will incur a latency overhead which becomes severe at large scale (e.g. 512 GPUs). FSDP alone will become less efficient due to the collective latency increasing linearly with the world size, resulting in FSDP collectives that cannot be hidden by the computation any more. To further scale out, one needs to consider combining model parallelism solutions such as TP and PP.
3.3.2 2D parallelism: Apply TP with FSDP
Model Parallelism (TP and PP) can help avoid the increased collective latency faced by scaling FSDP alone. TP can further lower the effective local batch size (to a minimum of 1TP $\frac{1}{\text{TP degree}}$ when the local batch size is set to 1), as TP sharded models on multiple GPUs work jointly to process the same batch. This can be vital for reducing peak memory usage so that training could fit in GPU memory (e.g. due to large model size or sequence length), or for strong scaling with fixed desired global batch size (e.g. due to training efficiency considerations).
In addition, TP performs feature dimension sharding. This can bring more optimized matrix multiplication shapes for better FLOP utilization.
As TP introduces extra blocking collectives, in practice TP is only applied within a node that have fast interconnect (NVLink). AsyncTP could improve the performance by fully overlapping the communication, but could not scale to multi-node easily, so the TP degree is usually limited to 8. This means to scale beyond an ultra-large number of GPUs (e.g. more than 4192 GPUs), we need 3D parallelism that combines PP.
3.3.3 3D Parallelism: Apply PP with 2D Parallelism
Compared to other model parallelisms, PP requires less communication bandwidth by virtue of only transmitting activations and gradients between stages in a P2P manner. It is especially useful (1) to further reduce FSDP communication latency when the FSDP world size becomes large again that FSDP+TP still exposes FSDP collectives; or (2) to train with bandwidth-limited cluster.
We note that, the performance of PP, and in particular the “bubble” size, could vary by pipeline schedules being used and the microbatch size, assuming fixed global batch size and the world size.
4 Demonstrating adaptability and extensibility
In this section, we demonstrate the adaptability and extensibility of TorchTitan by highlighting ongoing work and external contributions that showcase its ability to seamlessly integrate and compare new techniques, optimizations, and models.
4.1 Ongoing work: 4D parallelism and zero-bubble pipeline schedules
TorchTitan’s modular and extensible architecture enables the seamless integration of new techniques and optimizations. For instance, ongoing work includes incorporating Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023) to enable 4D parallelism, and leveraging the torch.distributed.pipelining package to support zero-bubble schedules (Qi et al., 2023). This demonstrates TorchTitan’s ability to adapt to evolving machine learning landscapes.
4.2 External contributions: Building and evaluating custom innovations
TorchTitan’s flexible architecture also empowers users to easily integrate and compare new innovations. By providing a modular and efficient test bed, TorchTitan enables users to rapidly benchmark new techniques, optimizations, and hardware on their training performance. This has already led to the refinement of a new production-grade dataloader, improvements in a new ZeRO implementation, advancements in an Adam-based optimizer, and the training of a top-tier diffusion model.
5 Related works
With the rapidly growing significance of LLMs  (Dubey et al., 2024; Achiam et al., 2023), there is substantial research and industry focus on improving infrastructure for training LLMs of various sizes. Since the very nature of these models are large, distributed training support becomes inevitable. Libraries like Megatron (Narayanan et al., 2021), DeepSpeed (Rasley et al., 2020), and PyTorch distributed (Pytorch native) (Paszke et al., 2019; Meta Platforms, Inc., 2024a) offer APIs to build distributed training workflows. NVIDIA NeMo (NVIDIA Corporation, 2024), built on Megatron-LM, offers a packaged solution for handling complex end-to-end model life-cycle from data curation to model deployment. Pytorch-native solutions like torchtune (Meta Platforms, Inc., 2024b) focus on fine-tuning LLMs in a simplified workflow. TorchTitan differs from these solutions by focusing on production grade pre-training using PyTorch-native APIs. The library is designed with elastic composability to accomodate the scale required to pre-train LLMs with minimal external dependencies. This lowers the bar to understand and extend pre-training, while offering features like async distributed checkpointing for building an end-to-end production workflow.
@@ 6 Conclusion
TorchTitan is a powerful and flexible framework for training LLMs. It offers composability, allowing users to combine various parallelism techniques (FSDP, TP, and PP), memory optimization methods (Float8 and activation checkpointing), and integration with torch.compile to optimize training efficiency. TorchTitan is highly flexible, adaptable to evolving model architectures and hardware advancements, and features a modular design with multi-axis metrics that foster innovation and experimentation. TorchTitan also prioritizes interpretability, production-grade training, and PyTorch native capabilities. Additionally, it provides high-performance training with elastic scalability, comprehensive training recipes and guidelines, and expert guidance on selecting and combining distributed training techniques. As shown in the experiment sections, TorchTitan provides training accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) over optimized baselines. With its robust features and high efficiency, TorchTitan is an ideal one-stop solution for challenging LLM training tasks.

### v3

@@ 2.2.1 Navigating compute-memory trade-offs using activation checkpointing
Activation checkpointing (AC) (Chen et al., 2016; He and Yu, 2023; Purandare et al., 2023) and selective activation checkpointing (SAC) (Korthikanti et al., 2023) are standard training techniques to reduce peak GPU memory usage, by trading activation recomputation during the backward pass for memory savings. It is often needed even after applying multi-dimensional parallelisms.
@@
Within a TransformerBlock, full AC works by recomputing all activation tensors needed during the backward pass, whereas op-level SAC saves the results from computation-intensive PyTorch operations and only recomputes others. Layer-level SAC works in similar fashion as full AC, but the wrapping is applied to every $x$ TransformerBlock (where $x$ is specified by the user) to implement configurable trade-offs between memory and recompute. (Details are in Appendix  8.6.)
@@
In TorchTitan, we use regional compilation, which applies torch.compile to each individual TransformerBlock in the Transformer model. This has two main benefits: (1) we get a full graph (without graph breaks) for each region, compatible with FSDP2 and TP (and more generally torch.Tensor subclasses such as DTensor) and other PyTorch distributed training techniques; (2) since the Llama model stacks identical TransformerBlock layers one after another, torch.compile can identify the same structure is being repeatedly compiled and only compile once, thus greatly reducing compilation time.
@@ 2.2.3 Asynchronous Tensor Parallel to maximally overlap communication
By default, TP incurs blocking communications before/after the sharded computations, causing computation resources to not be effectively utilized. Asynchronous TP (AsyncTP) (Wang et al., 2022) achieves computation-communication overlap by fractionalizing the TP matrix multiplications within attention and feed-forward modules into smaller chunks, and overlapping communication collectives in between each section. The overlap is achieved by a micro-pipelining optimization, where results are being communicated at the same time that the other chunks of the matmul are being computed.
PyTorch AsyncTP is based on a SymmetricMemory abstraction, which creates intra-node buffers to write faster communication collectives. This is done by allocating a shared memory buffer on each GPU in order to provide direct P2P access (PyTorch Team, 2024a).
@@ 2.2.4 Boosting throughput with mixed precision training and Float8 support
Mixed precision training (Micikevicius et al., 2018) provides both memory and computational savings while ensuring training stability. FSDP2 has built-in support for mixed precision training with basic torch.dtype. This covers the popular usage of performing FSDP all-gather and computation in a low precision (e.g. torch.bfloat16), and perform lossless FSDP reduce-scatter (gradient) in high precision (e.g. torch.float32) for better numerical results. See Appendix  8.8 for usage details.
TorchTitan also supports more advanced mixed precision training with Float8, a derived data type, applied selectively to linear layers (available on newer hardware like NVIDIA H100), achieving substantial performance gains while ensuring training stability (reported in Section 3.2). The Float8 feature from torchao.float8 supports multiple per-tensor scaling strategies, including dynamic, delayed, and static (see Micikevicius et al. (2022); PyTorch Community (2023), Section 4.3 for details), while being composable with other key PyTorch-native systems such as autograd, torch.compile, FSDP2 and TP (with Float8 all-gather capability) (PyTorch Team, 2024c).
@@ 2.3.1 Scalable and efficient Distributed Checkpointing
Checkpoint save/load are crucial in training large language models for two reasons: they facilitate model reuse in applications like inference and evaluation, and they provide a recovery mechanism in case of failures. An optimal checkpointing workflow should ensure ease of reuse across different parallelisms and maintain high performance without slowing down training. There are two typical checkpointing methods. The first aggregates the state (model parameters and optimizer states) into an unsharded version that is parallelism-agnostic, facilitating easy reuse but requiring expensive communication. The second method has each trainer save its local sharded state, which speeds up the process but complicates reuse due to embedded parallelism information.
DCP addresses these challenges using DTensor, which encapsulates both global and local tensor information independently of parallelism. DCP converts this information into an internal format for storage. During loading, DCP matches the stored shards with the current DTensor-based model parameters and optimizer states, fetching the necessary shard from storage. TorchTitan effectively uses DCP to balance efficiency and usability. Furthermore, DCP enhances efficiency through asynchronous checkpointing by processing storage persistence in a separate thread, allowing this operation to overlap with subsequent training iterations. TorchTitan utilizes DCP’s asynchronous checkpointing to reduce the checkpointing overhead by 5-15x compared to synchronous distributed checkpointing for the Llama 3.1 8B model (PyTorch Team, 2024b).
@@ 2.3.2 Flight Recorder to Debug Job Crashes
Debugging NCCL collective timeouts at large scales is challenging due to the asynchronous nature of communication kernels. PyTorch’s Flight Recorder addresses this by logging the start, end, and enqueue times for all collective and p2p operations, along with metadata like process groups, source/destination ranks, tensor sizes, and stack traces.
This data is invaluable for diagnosing hangs in parallelism code. For PP, it can pinpoint the latest send or recv completed on the GPU, helping debug schedule bugs. For FSDP and TP, it identifies ranks that failed to call collectives, aiding in uncovering issues with PP scheduling or TP logic.
@@ 3 Experimentation
In this section, we demonstrate the effectiveness of elastic distributed training using TorchTitan, via experiments on Llama 3.1 8B, 70B, and 405B, from 1D parallelism to 4D parallelism, at the scale from 8 GPUs to 512 GPUs. We also share the knowledge and experience gained through TorchTitan experimentation. A walkthrough of the codebase on how we apply (up to) 4D parallelism can be found in Appendix 7.
@@ 3.2 Performance
To showcase the elasticity and scalability of TorchTitan, we experiment on a wide range of GPU scales (from 8 to 512), as the underlying model size increases (8B, 70B, and 405B) with a varying number of parallelism dimensions (up to 4D). To demonstrate the effectiveness of the optimization techniques introduced in Section 2.2, we show how training throughput improves when adding each individual technique on appropriate baselines. In particular, when training on a higher dimensional parallelism with new features, the baseline is always updated to include all previous techniques.
@@
Table 1: 1D parallelism (FSDP) on Llama 3.1 8B model, 8 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 16. (Stats per GPU)
@@
Table 2: 1D parallelism (FSDP) on Llama 3.1 8B model, 128 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 256. (Stats per GPU)
@@
Table 3: 2D parallelism (FSDP + TP) + torch.compile + Float8 on Llama 3.1 70B model, 256 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 32, TP degree 8. Local batch size 16, global batch size 512. (Stats per GPU)
@@
Table 4: 3D parallelism (FSDP + TP + PP) + torch.compile + Float8 + AsyncTP on Llama 3.1 405B model, 512 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 4, TP degree 8, PP degree 16. Local batch size 32, global batch size 128. (Stats per GPU)
@@
Table 5: FSDP + CP + torch.compile + Float8 on Llama 3.1 8B model, 8 GPUs. Mixed precision training. Full activation checkpointing. Local batch size 1. (Stats per GPU)
Schedule Sequence Length Throughput (Tok/Sec) Memory (GiB)
FSDP 8, CP 1 32,768 3,890 83.9
FSDP 4, CP 2 65,536 2,540 84.2
FSDP 2, CP 4 131,072 1,071 84.0
FSDP 1, CP 8 262,144 548 84.5
Table 6: 4D parallelism (FSDP + TP + PP + CP) + torch.compile + Float8 + AsyncTP + 1F1B on Llama 3.1 405B model, 512 GPUs. Mixed precision training. Full activation checkpointing. TP degree 8, PP degree 8. Local batch size 8. (Stats per GPU)
Schedule Sequence Length Throughput (Tok/Sec) Memory (GiB)
FSDP 8, CP 1 32,768 76 75.3
FSDP 4, CP 2 65,536 47 75.9
FSDP 2, CP 4 131,072 31 77.1
FSDP 1, CP 8 262,144 16 84.9
Additional experimental details and loss-convergence tests for correctness can be found in Appendix 8.10.
3.3 Scaling with TorchTitan 4D Parallelism
Scaling large language models (LLMs) requires parallelism strategies to handle increasing model sizes and data on thousands of GPUs. TorchTitan enables efficient scaling through composable 4D parallelism. This section highlights key observations and motivations for using TorchTitan 4D parallelism, focusing on a specific combination shown in Figure 2.
Figure 2: Scaling with 4D Parallelism
@@ 3.3.1 Scaling with FSDP
FSDP (ZeRO) is a general technique applicable to any model architecture and is often sufficient as the first degree of parallelism when communication is faster than computation (e.g., up to 512 GPUs). However, with larger scales, collective latency increases linearly with the world size, limiting efficiency. To overcome this, model parallelism like TP and PP can be combined with FSDP.
3.3.2 2D Parallelism: TP with FSDP
Tensor Parallelism (TP) reduces collective latency by distributing work across GPUs, enabling smaller effective batch sizes and reducing peak memory usage for large models or sequence lengths. Combining FSDP and TP allows strong scaling with a fixed problem/batch size (Details shown in Figure 4). TP also improves FLOP utilization by optimizing matrix multiplication shapes. However, TP introduces blocking collectives and is typically limited to intra-node scaling (e.g., NVLink), with degrees usually capped at 8. Scaling beyond 4192 GPUs requires combining TP with PP.
3.3.3 3D Parallelism: PP with 2D Parallelism
Pipeline Parallelism (PP) reduces communication bandwidth requirements by transmitting only activations and gradients between stages in a peer-to-peer manner. PP is particularly effective for mitigating FSDP communication latency at larger scales or in bandwidth-limited clusters. The efficiency of PP depends on pipeline schedules and microbatch sizes, which influence the size of pipeline “bubbles.”
3.3.4 Long Context Training and 4D Parallelism
Context Parallelism (CP) allows ultra long context training by splitting the context (sequence) dimension across GPUs to avoid OOM errors. CP is mainly used for long context training, to give the model capability to capture more correlations for tokens, thus enhancing the overall model quality. For scaling sequence length, CP can be used alone or together with DP. When training large models or on large number of GPUs, we can combine CP with 3D paralleism, where TP usually keeps the innner-most DeviceMesh dimension, and CP applies in the next outer DeviceMesh dimension.
4 Related Work
Libraries such as Megatron-LM (Narayanan et al., 2021), DeepSpeed (Rasley et al., 2020), veScale (Inc., 2024) and PyTorch Distributed (Paszke et al., 2019; Meta Platforms, Inc., 2024) provide APIs for distributed workflows. However, these frameworks present challenges in flexibility, integration, and scalability. TorchTitan addresses these limitations with native support for key features absent in existing systems:
  • Megatron-LM: Requires model modifications for TransformerEngine, lacks seamless FSDP integration with TP and PP, and does not support advanced pipeline schedules to minimize computation overhead.
  • DeepSpeed: Depends on Megatron-LM for TP and CP, with limited support for FSDP and advanced pipeline schedules.
  • veScale: Does not support FSDP, CP, SAC, Float8 training, or torch.compile, and offers only three pipeline schedules, compared to TorchTitan ’s six.
We note that each of these libraries has its own strengths, and TorchTitan is designed to provide foundational components that can be leveraged by all of them. A detailed comparison, including feature breakdowns and code complexity analysis, is available in Appendix 8.9. Slapo (Chen et al., 2023) introduces a schedule language to convert a PyTorch model for common model training optimizations such as 3D parallelism, and supports progressive optimization through high-level primitives. In contrast, TorchTitan provides modular and composable APIs built on DTensor and DeviceMesh.
@@ 5 Conclusion
TorchTitan is a powerful and flexible framework for LLM training, enabling seamless composability of parallelism techniques (FSDP, TP, PP, CP), memory optimizations (Float8, activation checkpointing), and PyTorch compiler integration for enhanced efficiency. Its modular design supports evolving architectures and hardware, fostering innovation with multi-axis metrics.
Designed for interpretability and production-grade training, TorchTitan offers elastic scalability, comprehensive training recipes, and expert guidance on distributed training strategies. As demonstrated in experiments, it accelerates training by 65.08% on Llama 3.1 8B (128 GPUs, 1D), 12.59% on Llama 3.1 70B (256 GPUs, 2D), and 30% on Llama 3.1 405B (512 GPUs, 3D) over optimized baselines, while enabling long-context training with 4D composability. With its robust features and high efficiency, TorchTitan is an ideal one-stop solution for challenging LLM training tasks.